OpenStudy (anonymous):

find the solution of 7w-w^2=0

OpenStudy (anonymous):

\[7w-w^2=w(7-w)\]

OpenStudy (anonymous):

can you see what it is?

OpenStudy (anonymous):

factor out wtv u can so w(7-w)=0 so w=0 and 7-w=0 w=0, 7 :)

OpenStudy (anonymous):

Try not to give the answer so fast @Mythias